OpenStudy (annasam1813):

How do you simplify this expression?: -5+i/2i? I'm really confused by it. If somebody could help that would be great! :-)

OpenStudy (annasam1813):

The question didn't originally have brackets. :-) But I don't think the answer would change with or without the brackets would it?

OpenStudy (518nad):

(-5+i)/(2i) = -5/(2i) + i/(2i) = 5i/2 + 1/2

OpenStudy (518nad):

the answer would change without the brackets

OpenStudy (518nad):

it would be a completely real number without brackets

OpenStudy (amorfide):

You will multiply by 1. if you have  this is the same as  and 4/4 =1 so in your case, you would multiply by 2i/2i then simplyify